OpenStudy (bananas1234):

Simplify and select the answer with the appropriate restrictions for the variables. (x+4)(x+2)/4x+8

OpenStudy (bananas1234):

@JoannaBlackwelder

geerky42 (geerky42):

Factor 4 out in denominator (4x+8)

OpenStudy (bananas1234):

(x+4)(x+2)/(4x+8) = (x+4)(x+2)/4(x+2) = (x+4)/4 What do i do next?

OpenStudy (bananas1234):

@jim_thompson5910

OpenStudy (bananas1234):

my options are a : x+2/4, x is not = -2 b: x+4/4 x is not = -2 c: 4/x+4 , x is not = -4 d: 4/x+2 , x is not = -2

OpenStudy (bananas1234):

the = with the line though it is the same as is not right? because i wrote is not instead of the = with the line though.

geerky42 (geerky42):

There is only one option with (x+4)/4.

OpenStudy (bananas1234):

so b?

geerky42 (geerky42):

Yep..

OpenStudy (bananas1234):

thanks